[Brahms-l] Shift report 20040429 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Thu Apr 29 2004 - 07:32:23 EDT
------------------------------------------------------------
Date: Thu Apr 29 07:32:22 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 116 - 116

Shift Summary:
------------------------------------------------------------
1.  We had beam when I started the shift.  They had problems trying to refill RHIC, but around 4:30 they declared physics running.

2. I continued taking data for FS at 4 deg at 1/1 A. 

3. I notice no detector problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 29 07:32:35 2004

This archive was generated by hypermail 2.1.8 : Thu Apr 29 2004 - 07:32:57 EDT